# Explicit estimates for the logarithmic derivative and the reciprocal of the Riemann zeta-function

### v3

@@ 5.1 Assuming a partial Riemann hypothesis
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\log\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|+\int_{1-\delta_{t}}^{1+\delta_{1,t}}\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(x+it)\right|\text{d}x,$ (47)
where the integral is bounded from above by
$\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right).$
To bound the first term on the right-hand side of (47), we use Lemma 12 with $\kappa=\delta_{1,t}=d_{1}/\log t$ to see that
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|\leq V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$ (48)
Hence (47) becomes
$\left|\frac{1}{\zeta(\sigma+it)}\right|$ $\leq\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}V\left(\frac{d_{1}}{\log t_{0}},\sigma_{1},t_{0},\eta\right)\log t,$
for $t_{0}\leq t\leq T-1/2$ and $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
Now we turn to the case $\sigma\geq 1+\delta_{1,t}$. Here we use the trivial bound
$|\zeta(s)|\leq\zeta(\sigma)\leq\frac{\sigma}{\sigma-1},\qquad(\sigma>1),$ (49)
where $\sigma/(\sigma-1)$ is decreasing in $\sigma$. Together with the trigonometric inequality from (25), we see that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\zeta(1+\delta_{1,t})\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t.$ (50)
We reiterate that although there are sharper bounds available to use here, this is sufficient for our purposes as the case where $\sigma\leq 1+\delta_{1,t}$ is dominant when computing constants.
Note that the results proved hold also if one assumes $RH$, and allow us to remove the upper bound restriction on $t$, as explained at the end of the proof of Theorem 2. ∎
5.2 Above the Riemann height
Now we consider $t\geq H$ and prove the following.
Theorem 5.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d_{1}$, $\sigma_{1}$, $W$, $d$, such that $W=1/d$, and $1+d_{1}/\log H\leq\sigma_{1}\leq 2$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y_{H}(d,d_{1},\sigma_{1},\eta)\log t,$
where
$Y_{H}$ $(d,d_{1},\sigma_{1},\eta)=\max\Bigg{\{}\frac{1}{d_{1}}+\frac{1}{\log H},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}V\left(\frac{d_{1}}{\log H},\sigma_{1},H,\eta\right)\Bigg{\}},$
with
$A_{5}$ $=\left(1+\frac{R^{\prime}}{H}\right)^{1/6}\left(1+\frac{\log\left(1+\frac{R^{\prime}}{H}\right)}{\log H}\right)\left(1+a_{0}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)$
×$\qquad\qquad\bm{\times}0.618\left(1+a_{1}\left(\frac{3}{2}+\frac{d_{1}-d}{\log H},1.31,H-R^{\prime}\right)\right)^{7/6},$
and $V(d_{1}/\log H,\sigma_{1},H,\eta),\,\eta$ are defined in Lemma 12. Admissible values of $Y_{H}$ with corresponding $W$ are presented in Table 4.
Proof.
The proof is analogous to that of Theorem 4. Let $\delta_{t}:=d/\log t$ and $\delta_{1,t}:=d_{1}/\log t$, and first suppose $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$. As we do not assume $RH(T)$, we have to check that (3) holds, and we do this by our hypothesis of $d$ satisfying (37). The same argument in Section 4.2 with Lemma 2 leading to (42), requires a uniform bound on $|\zeta(s)|$ in the region
$\frac{1}{2}\leq 1-\delta_{t}-R^{\prime}\leq\sigma\leq 1+\delta_{1,t}+R^{\prime}\leq\frac{3}{2}+\frac{d_{1}-d}{\log H},$
and in return gives us
@@ 5.2 Above the Riemann height
$\left|{\textup{Re}}\frac{\zeta^{\prime}}{\zeta}(\sigma+it)\right|\leq\frac{4}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\log t$
uniformly for $1-\delta_{t}\leq\sigma\leq 1+\delta_{1,t}$.
The rest of the proof then follows that of Theorem 4, and it remains to optimise $Y_{H}$ for a fixed choice of $W=1/d$, over $d_{1}$, $\sigma_{1}$ and $\eta$, subject to the conditions mentioned in the statement of the theorem, while also checking that (11) holds. ∎
$W$ $Y_{H}$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.5587$ $2.307$ $1.059544$ $1.412209$ $3.464102$
$6$ $2.284$ $1.060503$ $1.608124$ $3.775135$
$7$ $2.244$ $1.062065$ $1.604395$ $3.772048$
$8$ $2.214$ $1.063547$ $1.895304$ $3.621918$
$9$ $2.191$ $1.064570$ $1.399356$ $3.464227$
$10$ $2.173$ $1.065554$ $1.135564$ $3.65928$
$11$ $2.158$ $1.065999$ $1.43905$ $3.446264$
$12$ $2.146$ $1.066605$ $1.345521$ $3.769427$
$13$ $2.136$ $1.067103$ $1.345557$ $3.77319$
Table 4: Values for $Y_{H}$ in Theorem 5, with corresponding $W$, and parameters (rounded to six decimal places) $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq H$.
5.3 Combining both cases $t\leq H$ and $t\geq H$
@@
For the first assertion of the corollary, we begin by assuming $RH(H_{0})$ and dealing with the case $t_{0}\leq t\leq H$.
To obtain $Y(W,5)$, fix $t_{0}$ and $W=1/d$ in Theorem 4 and optimise $Y_{RH(H_{0})}$ over $d_{1}$, $\sigma_{1}$, and $\eta$. At the same time ensure that the restrictions set in Theorem 4 and Corollary 8 hold, and raise $t_{0}$ if (11) is not satisfied. We find that the smallest value of $t_{0}$ such that (11) is fulfilled for all entries in Tables 6 and 5 is $t_{0}=5$. Then we compare the resulting values with the corresponding $Y_{H}$ (for $t\geq H$) in Theorem 5 and Table 4, taking the larger of the two.
[…]
For the second assertion, i.e., when $\sigma=1$, we set $d=0$ and repeat the above process using Theorems 4 (under $RH(H_{0})$) and 5 (unconditionally). This allows us to obtain a result for $\sigma\geq 1$. We have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.018\log t\quad\text{for}\quad t\geq H,$ (51)
[…]
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq 2.751\log t\quad\text{for}\quad 2\exp(e^{2})\leq t\leq H,$ (52)
[…]
Finally, combine (51) and (52) with the following two bounds. First, we utilise Lemma 7, which was proven using numerical methods. Taking $\sigma=1$ in said lemma, we derive
$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
@@
Second, a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500.$
This completes the proof. ∎
When $\sigma$ is fixed (in this case $\sigma=1$), the limiting factor for the constant in the second assertion of Corollary 4 comes from the case $2\exp(e^{2})\leq t\leq H$. In particular, the lower threshold of $2\exp(e^{2})$ is what determines the constant $2.751$. Using numerical methods, it is possible to push this threshold higher and hence improve (52). To illustrate this idea, following the proof of Corollary 4, we can prove unconditionally that for $t\geq 10^{6}$,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.391\log t,$ (53)
which is an improvement over what one gets from [5, Theorem 1.1].
We make the remark that we include our computed values of $\sigma_{1}$ and $\eta$ in Tables 4, 5, and 6, for completeness sake. In practice, as we limit the overall constant to three decimal places, $d_{1}$ is the main parameter we optimise over. For $\sigma_{1}$ and $\eta$, any sufficiently large values will do, and further optimisations here will only improve the final constant from the fourth decimal place onwards.
$t_{0}$ $Y(5.5587,t_{0})$ $d_{1}$ $\sigma_{1}$ $\eta$
$10$ $8.101$ $0.307606$ $1.330692$ $3.068541$
$10^{2}$ $4.339$ $0.542825$ $1.851563$ $3.417699$
$10^{3}$ $3.632$ $0.648814$ $1.282661$ $3.465621$
$10^{4}$ $3.264$ $0.725026$ $1.748467$ $3.463629$
$10^{5}$ $3.021$ $0.787049$ $1.208003$ $3.464054$
$10^{7}$ $2.711$ $0.885011$ $1.101137$ $3.464109$
$10^{9}$ $2.518$ $0.960512$ $1.272668$ $3.445896$
$10^{11}$ $2.383$ $1.021396$ $1.655004$ $3.464097$
@@
Table 5: Values for $Y(5.5587,t_{0})$ in Corollary 4, with corresponding $t_{0}$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}$ and $W=5.5587$.
$W$ $Y(W,5)$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.5587$ $18.060$ $0.170031$ $1.105647$ $2.784753$
$6$ $16.722$ $0.173040$ $1.107519$ $2.781151$
$7$ $14.654$ $0.178370$ $1.110828$ $2.783844$
$8$ $13.339$ $0.182334$ $1.113292$ $2.786533$
$9$ $12.433$ $0.185403$ $1.115198$ $2.782321$
$10$ $11.771$ $0.187839$ $1.116712$ $2.782458$
$11$ $11.268$ $0.189804$ $1.117933$ $2.786745$
$12$ $10.872$ $0.191448$ $1.118954$ $2.786457$
$13$ $10.553$ $0.192838$ $1.119819$ $2.786611$
Table 6: Values for $Y(W,5)$ in Corollary 4, with corresponding $W$, and parameters for the case $t_{0}\leq t\leq H$ (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq t_{0}=5$.
@@ 5.4 An asymptotic improvement
In this section we prove an asymptotic improvement over the bounds given earlier in Section 5, at the expense of a worse constant. For the sake of clarity, we first outline the general idea here, before proving the next two theorems.
Assume we have the bound
@@
and at the expense of a worse constant, we can estimate the second factor in (55) differently from the previous sections. We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$. Along with Lemma 9, this leads to
@@
Assume $RH(T)$. Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log t_{0}\leq 1$. Suppose also $R^{\prime}=1/2-d/\log t_{0}$. Then in the region
$\sigma\geq 1-\frac{1}{W\log t}\qquad\text{and}\qquad R^{\prime}+e\leq t_{0}\leq t\leq T-\frac{1}{2},$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{RH(T)}(d,d_{1},t_{0})(\log t)^{11/12},$
@@
$Y^{\prime}_{RH(T)}$ $(d,d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+B_{t_{0}}+\frac{\log A_{4}}{\log t_{0}}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{4}$ defined in Theorem 4, $B_{t_{0}}$ defined in Theorem 2, and $C_{3}$ defined in Lemma 6.
@@
The proof is identical to that of Theorem 4, except that for $\sigma\leq 1+\delta_{1,t}$, by Lemmas 6 and 9, we can replace (48) with
$\left|\frac{1}{\zeta(1+\delta_{1,t}+it)}\right|$ $\leq|\zeta(1+\delta_{1,t})|^{3/4}|\zeta(1+\delta_{1,t}+2it)|^{1/4}$
$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}.$
Meanwhile, for $\sigma\geq 1+\delta_{1,t}$, the same argument with (24) and (49) leads to
@@
Theorem 7.
Let $\gamma$ be Euler’s constant and let there be positive parameters $d$, $d_{1}$, $W$, such that $W=1/d$, and $d_{1}/\log H\leq 1$. Suppose also $R^{\prime}=1/2-d/\log H$, and $W>W_{0}$ is chosen such that $d$ satisfies (37). Then in the region
$\sigma\geq 1-\frac{1}{W\log t},\qquad t\geq H,$
we have
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{H}(d,d_{1})(\log t)^{11/12},$
where
$Y^{\prime}_{H}$ $(d,d_{1})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log H}\right)^{3/4},$
$\exp\Bigg{(}\frac{4(d_{1}+d)}{\pi R^{\prime}}\left(\frac{1}{6}+\frac{\log\log H}{\log H}+\frac{\log A_{5}}{\log H}\right)\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log H}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $A_{5}$ defined in Theorem 5, and $C_{3}$ defined in Lemma 6.
Proof.
The proof is similar to that of Theorem 6, except we do not assume $RH(T)$. Then the process follows the proof of Theorem 5, making similar changes as with the earlier case $t\leq H$ in Theorem 6. ∎
Combining Theorems 6 and 7, we prove Corollary 5, a result which holds for all $t\geq t_{0}\geq 5$.
@@
The proof is analogous to that of Corollary 4, where the final result is obtained by optimising over the appropriate parameters.
For the result on $\sigma=1$, the limiting region is again $2\exp(e^{2})\leq t\leq H$, where we have
$\left|\frac{1}{\zeta(1+it)}\right|\leq 7.475(\log t)^{11/12},$
with $d_{1}=0.72042$. ∎
$t_{0}$ $Y^{\prime}(5.5587,t_{0})$ $d_{1}$ $W$ $Y^{\prime}(W,5)$ $d_{1}$
$10$ $21.123$ $0.307588$ $5.5587$ $47.054$ $0.170052$
$10^{2}$ $11.643$ $0.542830$ $6$ $43.568$ $0.173052$
$10^{3}$ $9.842$ $0.648792$ $7$ $38.179$ $0.178401$
$10^{4}$ $8.885$ $0.725028$ $8$ $34.753$ $0.182368$
$10^{5}$ $8.247$ $0.787048$ $9$ $32.391$ $0.185428$
$10^{7}$ $7.425$ $0.885011$ $10$ $30.668$ $0.187861$
$10^{9}$ $6.907$ $0.960531$ $11$ $29.356$ $0.18984$
$10^{11}$ $6.545$ $1.021397$ $12$ $28.325$ $0.191481$
$H$ $6.340$ $1.059544$ $13$ $27.494$ $0.192568$
Table 7: Values for $Y^{\prime}(W,t_{0})$ in Corollary 5, with corresponding $W$ or $t_{0}$, and parameter $d_{1}$ for the case $t_{0}\leq t\leq H$ (rounded to six decimal places). Each entry is valid for $t\geq t_{0}$.

### v4

@@ 4.3 Combining both cases $t\leq H$ and $t\geq H$
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|$ $=-\textup{Re}\log\zeta\left(1+\delta_{1}+it\right)+\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (44)
Writing $\Delta=1/\log t$, we can split the integral and rewrite it as
@@
$\left(\int_{1-\tfrac{\Delta}{W_{1}}}^{1-\tfrac{\Delta}{W_{2}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{j}}}^{1-\tfrac{\Delta}{W_{j+1}}}+\cdots+\int_{1-\tfrac{\Delta}{W_{J}}}^{1}+\int_{1}^{1+\delta_{1}}\right)\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x.$ (45)
By Corollary 2 and our assumptions we have
$\int_{\sigma}^{1+\delta_{1}}\textup{Re}\left(\frac{\zeta^{\prime}}{\zeta}(x+it)\right)\text{d}x$ $\leq\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
We therefore have, from (44),
$\log\left|\frac{1}{\zeta(\sigma+it)}\right|<$ $-\log\left|\zeta\left(1+\frac{d_{1}}{\log t}+it\right)\right|$ (46)
$+\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}.$
To estimate the first term on the right-hand side of (46) we apply Lemma 12, keeping in mind our restrictions on $\sigma_{1}$. This gives us
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|\leq V(d_{1}/\log t_{0},\sigma_{1},t_{0},\eta)\log t.$ (47)
We combine this with (46) and exponentiate to obtain $Y_{0}$ for $\sigma\leq 1+\delta_{1}$.
For the case $\sigma>1+\delta_{1}$ we have from the trivial bound $|1/\zeta(s)|\leq\zeta(\sigma)$ with (9) that
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)\log t,$ (48)
since $\sigma/(\sigma-1)$ is decreasing in $\sigma$.
Taking the maximum from both cases completes the proof. ∎
Corollary 4 follows from Theorem 4 by computing $Y_{0}$ for specific choices of $t_{0}$ and $W_{j}$, and optimizing over $d_{1}$, $\sigma_{1}$, and $\eta$. These computed values of $Y_{0}$ are labelled $Y$, and are presented in Table 4, with its corresponding parameters (rounded to six decimal places). The process is the same as that of [7, Corollary 3], but for completeness sake, we repeat it here.
@@
After choosing $t_{0}=13$, we aim to minimise $Y_{0}$ by optimising over $d_{1}>0$, $\eta$, and $\sigma_{1}$, while making sure that the conditions on these parameters are fulfilled.
This optimisation process first requires fixing the desired $W_{1}$, which then tells us the number of $W_{j}$ we are summing over, since from Table 3 we always have $(W_{J},Q_{1,J})=(13,52.306)$. Computing more values of $(W_{J},Q_{1,J})$ will give a better result, but the improvements eventually become negligible, as $Y_{0}$ is largely determined by the initial few $W_{j}$. For our computations, we only used values for $W_{j}=W$ from Table 3 (Corollary 1), thus we require $t_{0}=13$.
Note that one can adjust Theorem 4 to compute $Y$ for the case $\sigma\geq 1$. In that case, we will modify (43) by setting
$\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}=0$
before optimising. We obtain $Y=30.812$ for $\sigma\geq 1$ and $t\geq 13$, with $(d_{1},\sigma_{1},\eta)=(0.032871,1.662479,3.216997)$. This proves the second assertion of the corollary.
To prove the third assertion, we first compute $Y=29.388$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $(d_{1},\sigma_{1},\eta)=(0.034172,1.286253,3.718165)$. Then see that from Lemma 7,
$\left|\frac{1}{\zeta(1+it)}\right|\leq\frac{8}{\log(500)}\log t\leq 1.288\log t\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}),$
@@
while a computation from the proof of [3, Proposition A.2], implemented in interval arithmetic, states that
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.079\log t\quad\text{for}\quad 2\leq t\leq 500,$ (49)
so we are done. ∎
$W$ $Y$ $d_{1}$ $\sigma_{1}$ $\eta$
$5.56$ $7.59\cdot 10^{31}$ $0.030648$ $1.151072$ $3.150608$
66 $2.46\cdot 10^{9}$ $0.030648$ $1.150464$ $3.150009$
77 $1.83\cdot 10^{6}$ $0.030647$ $1.151694$ $3.149136$
88 $89357$ $0.030647$ $1.148849$ $3.149156$
99 $19482$ $0.030647$ $1.150203$ $3.150832$
$10$ $7589$ $0.030648$ $1.149868$ $3.150556$
$11$ $3972$ $0.030650$ $1.150918$ $3.151137$
$12$ $2472$ $0.030646$ $1.148656$ $3.150041$
$13$ $1719$ $0.030647$ $1.149567$ $3.150198$
@@
Table 4: Values for $Y$ in Corollary 4, with corresponding $W$, and parameters (rounded to six decimal places): $d_{1}$, $\sigma_{1}$, and $\eta$. Each entry is valid for $t\geq 13$.
@@ 5.1 An asymptotic improvement
In this section we prove an asymptotic improvement over the bounds given earlier in §5, at the expense of a worse constant. For the sake of clarity, we first outline the general idea here, before proving the next theorem.
Suppose we have the bound
@@
and at the expense of a worse constant, we can estimate the second factor in (51) differently from Theorem 4. We do this by using Lemma 6, which states that $|\zeta(s)|\leq C_{3}(\log t)^{2/3}$ for $\sigma\geq 1$. Along with the second inequality in Lemma 9 and our choice of $\delta_{1}$, this leads to
@@
Let the hypotheses of Theorem 4 be satisfied. Then for $t\geq t_{0}\geq 13$ and
$\sigma\geq 1-\frac{1}{W_{1}\log t},$
@@
$\left|\frac{1}{\zeta(\sigma+it)}\right|\leq Y^{\prime}_{0}(d_{1},t_{0})(\log t)^{11/12},$
@@
$Y^{\prime}_{0}$ $(d_{1},t_{0})=C_{3}^{1/4}\cdot\max\Bigg{\{}\left(\frac{1}{d_{1}}+\frac{1}{\log t_{0}}\right)^{3/4},$
$\exp\Bigg{(}\sum_{j=1}^{J-1}Q_{1,j}\left(\frac{1}{W_{j}}-\frac{1}{W_{j+1}}\right)+\frac{Q_{1,J}}{W_{J}}+24.303d_{1}\Bigg{)}\left(\frac{e^{\gamma d_{1}/\log t_{0}}}{d_{1}}\right)^{3/4}\Bigg{\}},$
with $d_{1}$, $J$, $Q_{1,j}$, $W_{j}$ defined as in Theorem 4, and $C_{3}$ defined as in Lemma 6.
@@
The proof is identical to that of Theorem 4, except that in the case $\sigma\leq 1+\delta_{1}$, we replace (47) with
$\left|\frac{1}{\zeta(1+\delta_{1}+it)}\right|$ $\leq|\zeta(1+\delta_{1})|^{3/4}|\zeta(1+\delta_{1}+2it)|^{1/4}$
$\leq\left(\frac{e^{\gamma(d_{1}/\log t_{0})}}{d_{1}}\log t\right)^{3/4}\left(C_{3}(\log t)^{2/3}\right)^{1/4}$
by using (51) with Lemmas 6 and 9.
Meanwhile, for $\sigma>1+\delta_{1}$, the same argument with (20) and (9) leads to
@@
The proof, and process of optimisation here, even for the case $\sigma\geq 1$, are exactly the same as that of Corollary 4. These computed values of $Y^{\prime}_{0}$ for $t\geq 13$, are labelled $Y^{\prime}$, and are presented in Table 5. For all these cases, we found an optimal parameter of $d_{1}=0.030648$ (rounded to six decimal places).
For the case $\sigma\geq 1$, we computed $Y^{\prime}=87.725$ for $t\geq 13$, also with optimal $d_{1}=0.030648$ (rounded).
Finally, for the case $\sigma=1$, we computed $Y^{\prime}=81.900$ for $\sigma\geq 1$ and $t\geq 2\exp(e^{2})$, with $d_{1}=0.030793$ (rounded). Then by a similar calculation as in the proof of Corollary 4, we have by Lemma 7,
$\left|\frac{1}{\zeta(1+it)}\right|\leq 1.5(\log t)^{11/12}\quad\text{for}\quad 500\leq t\leq 2\exp(e^{2}).$
On the other hand, (49) implies
$\left|\frac{1}{\zeta(1+it)}\right|\leq 2.421(\log t)^{11/12}\quad\text{for}\quad 2\leq t\leq 500,$
where this estimate is obtained by noting that $t$ attains its maximum at $500$. ∎
$W$ $Y^{\prime}$
$5.56$ $2.17\cdot 10^{32}$
66 $6.99\cdot 10^{9}$
77 $5.21\cdot 10^{6}$
88 $2.55\cdot 10^{5}$
99 $55579$
$10$ $21649$
$11$ $11331$
$12$ $7051$
$13$ $4904$
Table 5: Values for $Y^{\prime}$ in Corollary 5, with corresponding $W$. Each entry is valid for $t\geq 13$.